A 1,000 milliliter vial of medication is to be used for 3 patients. Joe is to receive half as much as Jim and 100 milliliters less than Jack. How much will each patient be administered?
Answer by stanbon(75887)

A 1,000 milliliter vial of medication is to be used for 3 patients. Joe is to receive half as much as Jim and 100 milliliters less than Jack. How much will each patient be administered?
-----
O + I + A = 1000 ml
O = (1/2)I
O = A - 100
------------------------
Substitue for "I" and for "A" and solve for "O"::
O + 2*O + O+100 = 1000
4*O = 900
O = 225 ml
I = 2*O = 450 ml
A = O+100 = 325 ml
--------------
Cheers,
Stan H.
------------